OpenStudy (anonymous):

collin left the party one half hour before gabriel. terrence left an hour and a half later than gabriel. together the three of them stayed more than four hours. how long did they each stay at the party?

OpenStudy (anonymous):

Are you sure you type the question correctly? It is too ambiguous to have one answer.

OpenStudy (anonymous):

oh i forgot a part to translate into inequality?

OpenStudy (anonymous):

c + g + t = 4 g + .5 = c t + 1.5 = g you can solve those with substitution

OpenStudy (anonymous):

It says they all stayed together more than 4 hours. What if Colin stayed for 20 hours, terrance Stayed for 22 hours, and gabriel stayed for 20.5 hours? That is a valid answer, the way this question is typed. Did you mean to say that they all stayed PRECISELY four hours at the party?

OpenStudy (anonymous):

No, it can be at least 4 hours. You can find the minimum times with the equations i posted before. The only precise thing you would have to change is \[c + g + t \ge 4\] You can still solve with inequalities.

OpenStudy (anonymous):

So you're saying: c + g + t >=4 c = g - .5 t = g + 1.5 So we have g - .5 + g + g + 1.5 >= 4 3g + 1 >= 4 3g >= 3 g >= 1 Right?

OpenStudy (anonymous):

The idea is correct, but you substituted t incorrectly. t + 1.5 = g not -1.5. Other than that, once you have g, you can solve c and t. I was going to let salvarican do that math.

OpenStudy (anonymous):

terrance left an hour and a half later than Gabriel, this implies T = G + 1.5 ?

OpenStudy (anonymous):

Oh yeah, my bad, copied the equation down wrong. Good catch!

OpenStudy (anonymous):

I think whoever formulated the problem runs a scabby operation anyway

OpenStudy (anonymous):

I've actually seen this question on here multiple times. It must be from a common algebra book.

OpenStudy (anonymous):

scabby operation i tells ya